undeadknight26 (undeadknight26):

Help please? I dunno where to star D:

undeadknight26 (undeadknight26):

1. To open these doors, you must match the number and type of solutions for the following two functions in standard form. o f(x) = x2 + 6x -16 o g(x) = x2 +6x + 1

OpenStudy (gorv):

solution means roots of the equation

OpenStudy (gorv):

u know how to find the root??

OpenStudy (gorv):

x^2+6x+16=0 can u break the middle term such that id broken part added they give =6 and their product =16

OpenStudy (gorv):

@undeadknight26

undeadknight26 (undeadknight26):

ummm o.o I dun remember learning this o.o

OpenStudy (gorv):

factorise 16

OpenStudy (gorv):

2*2*2*2

OpenStudy (gorv):

pick two factor here u see that 8-2=6

OpenStudy (gorv):

x^2+8x-2x+16=0

undeadknight26 (undeadknight26):

I see so we are like expanding the equation?

OpenStudy (gorv):

@undeadknight26 can u proceed it further??

OpenStudy (gorv):

yeah now take x common fromfirst 2 term and -2 from last 2 terms..n tell what u got??

undeadknight26 (undeadknight26):

Would you expand 16 also?

OpenStudy (gorv):

no no we need to find roots of the equation it is a qudratic equation

OpenStudy (gorv):

x(x+8)-2(x+8) (x+8)(x-2) so solution are x=2 and x=-8

OpenStudy (gorv):

similarly try other one

undeadknight26 (undeadknight26):

@gorv i got: g(x) = x2 +6x + 1 x^2 + x8 – 2x + 1 x(x + 8) -2x + 1

OpenStudy (gorv):

no buddy productof splitted part should be equal to 1

OpenStudy (gorv):

weneed to use quadratic formula

OpenStudy (gorv):

\[x=\frac{ -b \pm \sqrt{b^2-4*a*c}}{ 2*a }\]

OpenStudy (gorv):

ax^2+bx+c compare with this x^2+6x+1 a=1, b=6 and c=1